A weight loss drug claims that taking their weight loss pills that you will loose weight. To test, 12 people were interview (see table of results below) before and aftertaking the drug. Test the claim at a 0.01 level of significance. 4. 1 2 3 4 5 6 7 89 10 11 12 Before 300 250 295 275 325 310 198 410 356625 256 312 After 175 295 300 210 175 190250 300 210 401 175 210
Answer #1

H_{0}:\mu _{d}=0

vs

H_{1}:\mu _{d}>0

We can use here,

t=\frac{\bar{d}}{\frac{S_{d}}{\sqrt{n}}}

=\frac{85.0833}{\frac{82.6773}{\sqrt{12}}}

=3.565

Degree of freedom =n-1 = 12 -1 = 11

P value =TDIST(3.565,11,1)...................using Excel command.

P value =0.0022 < 0.01

We have sufficient evidence to say that taking their weight loss pills that you will loose weight.

Numbor Betare t-Test: Paired Twsample tor Means 9120607 4219.1/1242 Medisi 11737,00リ1 AN/A 82·G77311 6035 53 Pecrzon Coreltio